OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y = x2 - x - 56. A. -56 B. 56 C. -225/4 D. 225/4 I got A

OpenStudy (luigi0210):

Find the derivative

OpenStudy (anonymous):

i use - B/2a right

OpenStudy (luigi0210):

D you know what a and b are?

OpenStudy (anonymous):

yeah equation is ax2+bx+c

OpenStudy (luigi0210):

Plug it in and show me what you get

OpenStudy (anonymous):

(1/2, 56.5)

OpenStudy (anonymous):

i got it luigi 

OpenStudy (luigi0210):

Might want to check again bud

OpenStudy (anonymous):

ohh  i did -(-1)/2(1)=1/2  then i did 1/2︿2-(1/2)-56

OpenStudy (luigi0210):

Yea, so what's : \[\Large \frac{1}{4}-\frac{1}{2}-56\]

OpenStudy (anonymous):

1/4 -1/2 =1/2

OpenStudy (anonymous):

then wont i do .5-56?

OpenStudy (luigi0210):

Convert it to decimals if that helps: 1/4=.25 1/2=.5 So what's .25-.5

OpenStudy (anonymous):

.20

OpenStudy (luigi0210):

Really? \[50-25=20?\]

OpenStudy (anonymous):

ohhh ok i see sorry .25-.50=.25

OpenStudy (anonymous):

i understod .05 XDmy bad

OpenStudy (luigi0210):

Yup, and since 50 is larger than 25, it is -.25 Now, what's (-.25)+(-56.00)=?

OpenStudy (anonymous):

-56.25

OpenStudy (luigi0210):

And there is your answer Mr Reaper

OpenStudy (luigi0210):

(1/2, -53.25)

OpenStudy (luigi0210):

*-56.25

OpenStudy (anonymous):

and what do i do i usually get stuck here

OpenStudy (luigi0210):

What do you mean?

OpenStudy (anonymous):

for the awnser what would i put a?

OpenStudy (luigi0210):

It'd be C, because -225/4=-56.25

OpenStudy (anonymous):

ohhh ok i see thanks :)

OpenStudy (luigi0210):

Yup!